# Demonstration of weighted graph optimization on a Rydberg atom array using local light-shifts

### v2

@@ Abstract
Neutral atom arrays have emerged as a versatile platform towards scalable quantum computation and optimization. In this paper we present first demonstrations of weighted graph optimization on a Rydberg atom array using annealing with local light-shifts. We verify the ability to prepare weighted graphs in 1D and 2D arrays, including embedding a five vertex non-unit disk graph using nine physical qubits and demonstration of a simple crossing gadget. We find common annealing ramps leading to preparation of the target ground state robustly over a substantial range of different graph weightings. This work provides a route to exploring large-scale optimization of non-planar weighted graphs relevant for solving relevant real-world problems.
@@
The ability to implement programmable spin models extends the utility of neutral atom arrays to the regime of analogue quantum computing [33, 34], where in the context of graph optimization coupling atoms to the Rydberg states enables native embedding of the Maximum Independent Set (MIS) problem for planar 2D unit-disk graphs (UDG) [35, 36]. Here the Rydberg blockade mechanism [37] prevents more than a single Rydberg excitation occuring within a radius $r<r_{\mathrm{B}}$, enforcing the constraint that no two connected vertices should be simultaneously part of the same MIS configuration. Experimental demonstrations of solving MIS on 2D-UDG using QAA with hundreds of qubits demonstrate the feasibility of this approach [38, 39], however recent analysis shows for the native Kings-graph classical solvers are able to find solutions to graphs with 1000 vertices in minutes [40]. One approach for extending beyond planar 2D graph is to use ancillary qubits to create quantum wires [41] which can be routed through neighboring trapping planes in 3D to implement non-local graph connectivity [42]. This approach has been used to demonstrate embedded integer factorization [43] and quadratic unconstrained binary optimization (QUBO) problems [44] onto atom arrays using only global control.
@@
Another approach to extend the range of problems that can be encoded using neutral atom arrays is to introduce local control fields as initially proposed in [45]. This permits the introduction of weighting to enable solution of Maximum Weighted Independent Set (MWIS) problems on a 2D UDG, providing a route to map a variety of graph and QUBO problems onto neutral atom arrays using this MWIS-UDG encoding [46] with at worst a quadratic overhead in physical qubit number. Similar scaling is obtained using a related parity-based encoding for graph problems [47]. Moving beyond the blockade regime to remove the UDG restriction offers prospects for a linear scaling in qubit overhead and application to MaxCut problems [45]. To date, local light-shifts have been used on neutral atom arrays in the field of quantum simulation for generating large-scale entangled states [48], implementing fast, arbitrary local control of interacting trimers [49] and studying quantum coarsening dynamics [50].
@@
In this paper we present first demonstrations of weighted graph optimization on a Rydberg atom array using annealing with local light-shifts. Our approach for weighting combines holographic techniques for generating programmable local light-shifts compatible with use on an arbitrary array geometry with the ability to implement graph annealing using global laser control fields, simplifying the number of available parameters when performing optimization. We verify the ability to prepare weighted graphs on both 1D and 2D arrays, including embedding a 5-qubit non-UDG using 9 physical qubits. In each case we find a common annealing ramp able to prepare the target ground state solutions for arbitrary graph weightings. This work demonstrates the feasibility of implementing weighted optimization using neutral atom arrays, and provides a route to embedding non-planar graphs solving relevant real-world problems [51].
@@
Weighted Graph Optimization.— The Maximum Weighted Independent Set (MWIS) problem on graph $G(V,E)$ comprized of $E$ edges and $V$ vertices, each of weight $w_{i}>0$, is defined as finding the subset of vertices offering the highest total summed weights whilst ensuring that no pair of adjacent vertices (i.e. connected by an edge) are included. Introducing the binary variables $n_{i}=\{0,1\}$ to indicate whether vertex $i$ belongs to the MWIS set, the classical cost function that must be minimized to solve the MWIS problem is
@@
For arbitrary problems, such as non-UDG graphs or QUBO, the graph (and associated constraints) must first be encoded into a weighted UDG-MWIS. Ref. [46] provides one approach to implementing this mapping by introducing regular sub-graph gadgets to realize arbitrary non-local connections between vertices, however the focus of this present work is to verify the experimental feasibility for preparing the MWIS on a neutral atom array.
@@
MWIS Annealing Protocol.— Starting with a target MWIS graph problem, the graph is first mapped to a UDG-MWIS which can be geometrically embedded onto a neutral atom array using techniques from [46] with a spacing $a$ between neighbouring atoms chosen to be below the blockade radius $r_{\mathrm{B}}$ to ensure the blockade constraint is realized. Graph weighting is implemented with an additional light-shifting potential that is holographically projected onto the atoms using a spatial light modulator (SLM) to apply local light-shifts with arbitrary geometry and programmable relative weighting. These additional tweezer beams, shown schematically in Fig. 1, cause a differential AC Stark shift $\delta_{\mathrm{AC}}^{i}$ on each atom proportional to the relative power in each tweezer, resulting in a total shift $\Delta_{i}=\Delta+\delta_{\mathrm{AC}}^{i}$, where $\Delta$ is the detuning of the global Rydberg lasers used to realize a homogeneous two-photon coupling between ground and Rydberg states. In the experiments below, the light-shift laser is blue-detuned for both ground and Rydberg states, ensuring that $\delta_{\mathrm{AC}}^{i}\geq 0$ as required. For a given graph weighting, the relative light-shifts are calibrated to realize $\delta_{\mathrm{AC}}^{i}=w_{i}\delta_{\mathrm{AC}}$, where $\delta_{\mathrm{AC}}$ is the global unit scale ($w=1$) of the light-shift and $w_{i}$ are the relative weights of the embedded graph problem. To avoid violating the blockade condition, it is necessary to constrain the maximum value of $\delta_{\mathrm{AC}}$ such that $\mathrm{max}_{i}(\delta_{\mathrm{AC}}^{i})<V(a)$, where $a$ is the interatomic spacing.
To adiabatically prepare atoms in the ground state encoding the solution of the UDG-MWIS problem, a cubic detuning profile is used  [52], as illustrated in Fig. 1(c). Starting with the atoms initially prepared in |$\ket{g}$, the global Rabi frequency $\Omega$ is ramped up using a fixed negative detuning after which the detuning is swept towards resonance. Once the global detuning reaches resonance, $\Delta=0$, the global detuning is held fixed and the remaining detuning profile is applied by dynamically varying the intensity of the light-shifting potential to control $\delta_{\mathrm{AC}}(t)$. By deriving the local light-shifts from a single laser, it is possible to use global intensity control for implementing arbitrary detuning ramps whilst maintaining fixed relative weightings on the graph. The Rabi frequency $\Omega$ is then ramped off whilst maintaining the light-shift, after which the encoded bitstrings $\{n_{i}\}$ are read out by fluorescence imaging. Atoms excited to the Rydberg state are ejected from the trap, whilst atoms in the ground state survive and appear in the images. Using a closed-loop optimization process, the cubic ramp parameters are adjusted to minimize the classical cost function $\langle H_{\mathrm{MWIS}}\rangle$ to maximize the probability of preparing the system in the MWIS ground state, as shown in Fig. 1(d). Following this optimization process, the solution of the target MWIS graph problem can then be determined from the most probable output bitstrings.
@@
Experimental Setup.— The experimental setup is based on arrays of holographically trapped Cs atoms. Atoms are initially prepared in the $|g\rangle=|4,0\rangle$ state, and are globally excited to the |$\ket{r}=\ket{80s_{1/2},m_{j}=+1/2}$ Rydberg state via two-photon excitation using the $7p_{1/2}$ intermediate state using light at 459 nm and 1039 nm. We operate with a detuning of $\Delta^{\prime}/2\pi=$502 MHz from the intermediate state, and achieve a Rydberg Rabi frequency of $\Omega/2\pi=2.70(2)$ MHz. Global detuning is controlled by adjusting the frequency using a double-passed acousto-optic modulator (AOM) in the 459 nm path. For $80s_{1/2}$ we calculate $C_{6}=-3376~{}\mathrm{GHz}\,\mu$m6 [53], resulting in a resonant blockade radius $r_{B}=\sqrt[6]{|C_{6}|/\Omega}=10.4~{}\mu$m.
@@
Local light-shifts are realized using light at 800 nm derived from a Ti:Sapph laser that is projected from a spatial light modulator (SLM) and overlapped with the main optical trapping array. To suppress mechanical forces coming from the repulsive optical potential, the light-shift spot arrays are re-imaged onto the atoms with a $1/e^{2}$ radius of $3.0(4)~{}\mu$m. For a given UDG-MWIS, the arrays of local light-shift beams are generated using a Gerchberg-Saxon algorithm with the desired relative weighting [54]. To verify the light-shifts experienced by the atoms, we perform spectroscopy on the |$\ket{g}\rightarrow\ket{50s_{1/2},m_{j}=+1/2}$ transition to suppress interactions, and perform a closed-loop feedback on the SLM potentials to achieve target weightings with $<2\%$ RMS errors in the relative weightings after 5 iterations  [52]. The resulting differential light-shift is calibrated as $\delta_{\mathrm{AC}}=0.28(2)$ MHz///mW, which is dominated by the ground state light-shift as the Rydberg AC shift calculated using the polarizability of a free electron corresponds to only 40 kHz [55].
@@
1D Weighted Graphs.— As an artificial first demonstration of the feasibility to perform weighted optimization using the dual-stage protocol defined above, we begin with a simple and well studied problem of a 1D weighted graph as illustrated in Fig. 2(a). For a uniformly weighted set of $N$ atoms with $w_{i}=1$ spaced such that $a<r_{\mathrm{B}}<2a$, for odd $N$ the ground state corresponds to a $Z_{2}$ ordered phase with Rydberg excitations localized on the odd sites as previously observed using global control only [28]. Introducing a relative weighting of $w_{i}=2$ on the even sites results in a modification of the MWIS ground state to prepare the inverted output string with Rydberg excitations now localized to the even sites.
[…]
2D Weighted Graphs.— Moving beyond the trivial 1D case, we consider the example of a non-UDG 2D MWIS shown in Fig. 3(a). This graph, featuring five vertices, has MWIS solutions corresponding to either qubits 1 and 3 being excited, or qubits 2, 4 and 5. To map this to a UDG-MWIS graph for implementation on our neutral atom array, we follow the protocol of Ref. [46] resulting in a 9-qubit graph requiring four ancilla qubits to implement the edges between vertices 1 and 2 and vertex 5. By symmetry this introduces two additional weightings, $w_{\alpha}$ and $w_{\beta}$ on the ancilla qubits. We use the weightings $w_{\alpha}=(w_{1}+w_{3})/2$ and $w_{\beta}=(w_{2}+w_{4}+w_{5})/2$ which guarantees that for any arbitrary weighting with $w_{i}>0$ we realize the MWIS solution as the ground state of the Rydberg interaction Hamiltonian.
Initially we consider the weighted graph instance indicated by the inset of Fig. 3(d) with $(w_{1},w_{2},w_{3},w_{4},w_{5})=(2,1,2,1,1)$ using a $3\times 3$ arrangement of atoms with spacing $a=8~{}\mu$m. As before, we perform closed-loop optimization of the annealing profile, resulting in a shorter annealing time of only 1.65 $\mu$s with the light-shift turning on after 0.6 $\mu$s, as shown in Fig. 3(b). For this optimization process we minimize $\langle H_{\mathrm{MWIS}}\rangle$ for the embedded 9 qubit graph (rather than simply just considering the non-embedded 5-qubit graph), as whilst the values of the ancilla qubits are not relevant to the problem solution, there are many possible states encoded in the 9-qubit bitstrings that feature the correct 5-vertex solution state but only one of which corresponds to the energetic ground state. This ensures the optimizer favours solutions that more closely follow adiabatic preparation of the ground states, which can be verified by observing the dynamical evolution in preparing the correct output bitstring. Fig. 3(c) shows the preparation probability as a function of time, revealing a smooth accumulation of population after 1.2 $\mu$s and no fast oscillations observed when exploiting shorter annealing profile durations. Following optimization of ramp parameters we obtain $\langle H_{\mathrm{MWIS}}\rangle=-4.0(1)$, compared to an idealized cost of -8.
@@
Gadgets.— As discussed above, a number of proposals exist to extend the range of target optimisation problems that can be solved using neutral atom arrays by converting to a UDG-MWIS [45, 46, 47]. One such approach introduces simple gadgets that provide a generalised framework for mapping problems to UDG-MWIS [46], and here we consider a proof-of-concept implementation of the crossing gadget using our local light-shift approach. This crossing gadget, shown in Fig. 5(a), implements an edge between qubits 1 and 2, and qubits 3 and 4. This requires four additional atoms, where the logical qubits (1-4) are assigned weight $w_{1}=1$, whilst the ancilla gadgets have $w_{2}=2.86$ to ensure the lowest energy state corresponds to an equal superposition of the four degenerate output configurations with either qubit 1 or 2 excited independent of 3 or 4. This deviates from $w_{2}=4$ presented in [46] due to the requirement to compensate for the effects of the $1/R^{6}$ interaction tails on the ancilla atoms. For these measurements the intermediate state detuning was increased to $\Delta^{\prime}/2\pi=1$ GHz and Rabi frequency reduced to $\Omega/2\pi=1.28$ MHz motivated by numerical modelling of the annealing process. As above, we then optimize the annealing ramp experimentally to minimize $\langle H_{\mathrm{MWIS}}\rangle$ for the problem resulting in the ramp shown in Fig. 5(b). The distribution of the final output strings are shown in Fig. 5(c), where the four target ground states are the most likely to be observed with approximately equal probability, and $\geq 20\%$ combined probability. Analysis of the next four most probable output strings show these correspond to the case of a target ground state with one of the logical qubit Rydberg excitation’s missing, likely due to limitations in the Rydberg detection fidelity. For embedding of larger graphs, improvements in the Rydberg state detection can be implemented using microwave ionization [56, 29], however these results verify the proof-of-concept approach for performing gadget encodings on the current architecture.

### v3

@@ Abstract
Neutral atom arrays have emerged as a versatile platform towards scalable quantum computation and optimization. In this paper we present demonstrations of solving maximum weighted independent set problems on a Rydberg atom array using annealing with local light-shifts. We verify the ability to prepare weighted graphs in 1D and 2D arrays, including embedding a five vertex non-unit disk graph using nine physical qubits and demonstration of a simple crossing gadget. We find common annealing ramps leading to preparation of the target ground state robustly over a substantial range of different graph weightings. This work provides a route to exploring large-scale optimization of non-planar weighted graphs relevant for solving relevant real-world problems.
I Introduction
@@
The ability to implement programmable spin models extends the utility of neutral atom arrays to the regime of analogue quantum computing [33, 34], where in the context of graph optimization coupling atoms to the Rydberg states enables native embedding of the Maximum Independent Set (MIS) problem for quasi-planar 2D unit-disk graphs (UDG) [35, 36]. A UDG is one in which only vertices within a given distance from eachtoher are connected by an edge, which can be naturally mapped to neutral atom systems using the Rydberg blockade mechanism [37] which prevents more than a single Rydberg excitation occurring within a radius $r<r_{\mathrm{B}}$. For MIS this enforces the constraint that no two connected vertices can be simultaneously part of the same set. Experimental demonstrations of solving MIS on 2D-UDG using QAA with hundreds of qubits demonstrate the feasibility of this approach [38, 39], however recent analysis shows for the native Kings-graph classical solvers are able to find solutions to graphs with 1000 vertices in minutes [40]. One approach for extending beyond planar 2D graph is to use ancillary qubits to create quantum wires [41] which can be routed through neighboring trapping planes in 3D to implement non-local graph connectivity [42]. This approach has been used to demonstrate embedded integer factorization [43] and quadratic unconstrained binary optimization (QUBO) problems [44] onto atom arrays using only global control.
@@
Another approach to extend the range of problems that can be encoded using neutral atom arrays is to introduce local control fields as initially proposed in [45]. This permits the introduction of vertex weighting to enable solution of Maximum Weighted Independent Set (MWIS) problems on a 2D UDG, providing a route to map a variety of graph and QUBO problems onto neutral atom arrays by using a UDG-MWIS encoding. Ref. [46] provides one such approach to implementing this as a general mapping by introducing sub-graph gadgets to realize arbitrary non-local connections between vertices, with at worst a quadratic overhead in physical qubit number. A similar $\mathcal{O}(N^{2})$ scaling is obtained using an alternative parity-based encoding for converted graph problems to UDG-MWIS [47]. Crucially this also enables a far more efficient encoding of arbitrary problems, such as non-UDG graphs or QUBO, by mapping to UDG-MWIS with overheads $\mathcal{O}(N^{2})$ than is the case for mapping to UDG-MIS with $\mathcal{O}(N^{6})$ [46]. Moving beyond the blockade regime to remove the UDG restriction offers prospects for a linear scaling in qubit overhead and application to MaxCut problems [45]. To date, local light-shifts have been used on neutral atom arrays in the field of quantum simulation for generating large-scale entangled states [48], implementing fast, arbitrary local control of interacting trimers [49] and studying quantum coarsening dynamics [50].
@@
In this paper we present demonstrations of solving maximum weighted independent set problems on a Rydberg atom array using annealing with local light-shifts. Our approach for weighting combines holographic techniques for generating programmable local light-shifts compatible with use on an arbitrary array geometry with the ability to implement graph annealing using a dual-stage annealing protocol with global laser control fields, simplifying the number of available parameters when performing optimization. We verify the ability to encode weighted graphs on both 1D and 2D arrays, including embedding a 5-qubit non-UDG using 9 physical qubits. In each case we find a common annealing ramp able to prepare the target ground state solutions for arbitrary graph weightings. This work demonstrates the feasibility of implementing weighted optimization using neutral atom arrays, and provides a route to embedding non-planar graphs solving a wide range of relevant real-world problems [51].
II Weighted Graph Optimization
@@
The Maximum Weighted Independent Set (MWIS) problem on graph $G(V,E)$ comprized of $E$ edges and $V$ vertices, with each vertex having weight $w_{i}>0$, is defined as finding the subset of vertices offering the highest total summed weights whilst ensuring that no pair of adjacent vertices (i.e. connected by an edge) are included. Introducing the binary variables $n_{i}=\{0,1\}$ to indicate whether vertex $i$ belongs to the MWIS set, the classical cost function that must be minimized to solve the MWIS problem is
@@
Solving MIS and MWIS on a UDG is known to be NP-hard [52]. State-of-the-art classical solvers for UDG-MWIS, a number of polynomial-time approximation schemes have been developed [53, 54], as well as heuristic algorithms based on iterative local-search [55] or memetic algorithms [56] which can be applied to find approximate solutions without any guarantee of optimality. For exact solutions, MWIS can be solved using branch-and-bound methods for small to intermediate scale graphs (typically hundreds of vertices), with graph reductions enabling efficient branch-and-reduce algorithms [57] for scaling to larger graph instances. Reduction rules for solving MWIS on sparse graphs were introduced in [58] who demonstrate the ability to solve sparse real-world graphs with up to a million vertices two orders of magnitude faster than exact local-search methods, with subsequent improvements providing an algorithm running time $\mathcal{O}(1.1443^{n})$ for graphs with an average degree of 3 or less [59]. However, for the UDG-MWIS embeddings realised on Rydberg atom systems the graph is not necessarily sparse, with union-jack like connectivity giving vertices with a maximum degree of 8 meaning we typically realise a dense unit disk graph for which the reductions introduced in [58] for sparse graphs are expected to offer limited advantage. Development of new approaches for solving UDG-MWIS using quantum hardware therefore provides an opportunity to solve large graphs without restrictions on sparsity.
Whilst for UDG-MIS there have been detailed studies benchmarking the scaling of state-of-the-art classical solvers against graphs with union-jack connectivity which show that classical optimal solutions can be found efficiently for graphs with thousands of vertices [40], a similar analysis for UDG-MWIS remains outstanding. In the results below we instead focus on experimental demonstrations of solving UDG-MWIS on neutral atom systems. We demonstrate the feasibility of the approach and outline techniques for extending this in future to tackle larger graphs in a regime competitive against classical algorithms.
III MWIS Annealing Protocol
Starting with a target MWIS graph problem, the graph is first mapped to a UDG-MWIS which can be geometrically embedded onto a neutral atom array using techniques from [46] with a spacing $a$ between neighbouring atoms chosen to be below the blockade radius $r_{\mathrm{B}}$ to ensure the blockade constraint is realized. Graph weighting is implemented with an additional light-shifting potential that is holographically projected onto the atoms using a spatial light modulator (SLM) to apply local light-shifts with arbitrary geometry and programmable relative weighting. These additional tweezer beams, shown schematically in Fig. 1, cause a differential AC Stark shift $\delta_{\mathrm{AC}}^{i}$ on each atom proportional to the relative power in each tweezer, resulting in a total shift $\Delta_{i}=\Delta+\delta_{\mathrm{AC}}^{i}$, where $\Delta$ is the detuning of the global Rydberg lasers used to realize a homogeneous two-photon coupling between ground and Rydberg states. In the experiments below, the light-shift laser is blue-detuned from the atomic transitions for both ground and Rydberg states, ensuring that $\delta_{\mathrm{AC}}^{i}\geq 0$ as required. For a given graph weighting, the relative light-shifts are calibrated to realize $\delta_{\mathrm{AC}}^{i}=w_{i}\delta_{\mathrm{AC}}$, where $\delta_{\mathrm{AC}}$ is the global unit scale ($w=1$) of the light-shift and $w_{i}$ are the relative weights of the embedded graph problem. To avoid violating the blockade condition, it is necessary to constrain the maximum value of $\delta_{\mathrm{AC}}$ such that $\mathrm{max}_{i}(\delta_{\mathrm{AC}}^{i})<V(a)$, where $a$ is the interatomic spacing.
To adiabatically prepare atoms in the ground state encoding the solution of the UDG-MWIS problem a dual-stage annealing protocol is used, as illustrated in Fig. 1(c). The detuning is controlled using a cubic profile (see App. B), motivated by recent experimental demonstrations [38] and features slow variation in detuning as it crosses resonance where the energy gap between ground and first excited states are minimised whilst changing faster at larger detunings where the gaps are typically larger.
Starting with the atoms initially prepared in |$\ket{g}$, the global Rabi frequency $\Omega$ is ramped up using a fixed negative detuning after which the detuning is swept towards resonance. Once the global detuning reaches resonance, $\Delta=0$, the global detuning is held fixed and the remaining detuning profile is applied by dynamically varying the intensity of the light-shifting potential to control $\delta_{\mathrm{AC}}(t)$. By deriving the local light-shifts from a single laser, it is possible to use global intensity control with an acousto-optic modulator for implementing arbitrary detuning ramps whilst maintaining fixed relative weightings on the graph. The Rabi frequency $\Omega$ is then ramped off whilst maintaining the light-shift, after which the encoded bitstrings $\{n_{i}\}$ are read out by fluorescence imaging. Atoms excited to the Rydberg state are ejected from the trap, whilst atoms in the ground state survive and appear in the images. Using a closed-loop optimization process, the cubic ramp parameters are adjusted to minimize the classical cost function $\langle H_{\mathrm{MWIS}}\rangle$ to maximize the probability of preparing the system in the MWIS ground state, as shown in Fig. 1(d). Following this optimization process, the solution of the target MWIS graph problem can then be determined from the most probable output bitstrings.
@@ III MWIS Annealing Protocol
IV Experimental Setup
@@
The experimental setup is based on arrays of holographically trapped Cs atoms. Atoms are initially prepared in the $|g\rangle=|4,0\rangle$ state, and are coherently driven to the |$\ket{r}=\ket{80s_{1/2},m_{j}=+1/2}$ Rydberg state by two-photon excitation via the $7p_{1/2}$ intermediate state using light at 459 nm and 1039 nm that globally illuminates the array. We operate with a detuning of $\Delta^{\prime}/2\pi=$502 MHz from the intermediate state, and achieve a Rydberg Rabi frequency of $\Omega/2\pi=2.70(2)$ MHz. Global detuning is controlled by adjusting the frequency using a double-passed acousto-optic modulator (AOM) in the 459 nm path. For $80s_{1/2}$ we calculate $C_{6}=-3376~{}\mathrm{GHz}\,\mu$m6 [60], resulting in a resonant blockade radius $r_{B}=\sqrt[6]{|C_{6}|/\Omega}=10.4~{}\mu$m. Annealing is performed with traps off to suppress losses from the tweezers.
@@
Local light-shifts are realized using light at 800 nm derived from a Ti:Sapph laser that is projected from a spatial light modulator (SLM) and overlapped with the main optical trapping array. The light-shift spot arrays are imaged onto the atoms with a $1/e^{2}$ radius of $3.0(4)~{}\mu$m (See App. A). This waist size is chosen as a balance between being large enough to suppress losses from the mechanical force of the repulsive optical potential, whilst reducing cross-talk between neighbouring sites. For a given UDG-MWIS, the arrays of local light-shift beams are generated using a Gerchberg-Saxon algorithm with the desired relative weighting [61]. To calibrate the light-shifts experienced by the atoms, we perform spectroscopy on the |$\ket{50s_{1/2},m_{j}=+1/2}$ Rydberg state where the lower lying state suppresses the effect of atom-atom interactions. We use the measured shifts to implement closed-loop feedback on the SLM potentials to achieve target weightings with $<2\%$ RMS errors in the relative weightings after 5 iterations. The resulting differential light-shift is calibrated as $\delta_{\mathrm{AC}}=0.28(2)$ MHz///mW, which is dominated by the ground state light-shift as the Rydberg AC shift calculated using the polarizability of a free electron corresponds to only 40 kHz [62].
V Results
V.1 1D Weighted Graphs
As an artificial first demonstration of the feasibility to perform weighted optimization using the dual-stage protocol defined in Sec. III, we begin with a simple and well studied problem of a 1D weighted graph as illustrated in Fig. 2(a). For a uniformly weighted set of $N$ atoms with $w_{i}=1$ spaced such that $a<r_{\mathrm{B}}<2a$, for odd $N$ the ground state corresponds to a $Z_{2}$ ordered phase with Rydberg excitations localized on the odd sites as previously observed using global control only [28]. Introducing a relative weighting of $w_{i}=2$ on the even sites results in a modification of the MWIS ground state to prepare the inverted output string with Rydberg excitations now localized to the even sites.
The resulting distributions of output strings are shown for both uniform and weighted graphs in Fig. 2(c,e) respectively. Using a common annealing profile, we observe the correct ground states as the most probable state outputs in both cases, with probabilities of $19(1)\%$ in both cases. The additional 9% peak in Fig. 2(e) with state index 93 corresponds to the target output state with the first qubit also lost. This appears higher than other states due to slightly increased single atom loss measured for this array site. Time evolution data is shown for each graph in Fig. 2(d,f) respectively, which clearly demonstrates the adiabatic nature of graph preparation at longer times and the emergence of the ground state ordering as expected.
V.2 2D Weighted Graphs
Moving beyond the trivial 1D case, we consider the example of a non-UDG 2D MWIS shown in Fig. 3(a). This graph, featuring five vertices, has MWIS solutions corresponding to either qubits 1 and 3 being excited, or qubits 2, 4 and 5. To map this to a UDG-MWIS graph for implementation on our neutral atom array, we follow the protocol of Ref. [46] resulting in a 9-qubit graph requiring four ancilla qubits to implement the edges between vertices 1 and 2 and vertex 5. By symmetry this introduces two additional weightings, $w_{\alpha}$ and $w_{\beta}$ on the ancilla qubits. We use the weightings $w_{\alpha}=(w_{1}+w_{3})/2$ and $w_{\beta}=(w_{2}+w_{4}+w_{5})/2$ which guarantees that for any arbitrary weighting with $w_{i}>0$ we realize the MWIS solution as the ground state of the Rydberg interaction Hamiltonian.
@@ V Results
Initially we consider the weighted graph instance indicated by the inset of Fig. 3(d) with $(w_{1},w_{2},w_{3},w_{4},w_{5})=(2,1,2,1,1)$ using a $3\times 3$ arrangement of atoms with spacing $a=8~{}\mu$m. As before, we perform closed-loop optimization of the annealing profile, resulting in a shorter annealing time of only 1.65 $\mu$s with the light-shift turning on after 0.6 $\mu$s, as shown in Fig. 3(b). For this optimization process we minimize $\langle H_{\mathrm{MWIS}}\rangle$ for the embedded 9 qubit graph (rather than simply just considering the non-embedded 5-qubit graph), as whilst the values of the ancilla qubits are not relevant to the problem solution, there are many possible states encoded in the 9-qubit bitstrings that feature the correct 5-vertex solution state but only one of which corresponds to the energetic ground state. This ensures the optimizer favours solutions that more closely follow adiabatic preparation of the ground states, which can be verified by observing the dynamical evolution in preparing the correct output bitstring. Fig. 3(c) shows the preparation probability as a function of time, revealing a smooth accumulation of population after 1.2 $\mu$s and no fast oscillations observed when exploiting shorter annealing profile durations. Following optimization of ramp parameters we obtain $\langle H_{\mathrm{MWIS}}\rangle=-4.0(1)$, compared to an idealized cost of -8. This discrepancy in the measured cost is simply a consequence of the finite state preparation and readout fidelities leading to a distribution of output strings, with all other strings making a more positive contribution to the average cost with zero weighting contributions from atoms in the ground state and strong positive energy penalties in the case of a blockade violation.
@@
V.3 Crossing Gadget